# Improvement of Spiking Neural Network with Bit Planes and Color Models

### v2

@@ Abstract
Spiking neural network (SNN) has emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint. However, their practical adoption is constrained by several challenges, prominently among them being performance optimization. In this study, we present a novel approach to enhance the performance of SNN for images through a new coding method that exploits bit plane representation. Our proposed technique is designed to improve the accuracy of SNN without increasing model size. Also, we investigate the impacts of color models of the proposed coding process. Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets. To the best of our knowledge, this is the first research that considers bit planes and color models in the context of SNN. By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more efficient and effective SNNs models in future researches and applications.
@@
Spiking neural network, spike coding, image classification
@@ I Introduction
SNNs utilize discrete spike events to transmit information between neurons, mirroring the asynchronous and event-driven communication observed in biological neurons[3]. This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[4]. The inherent sparsity and temporal coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [5].
Despite their potential, SNNs pose significant challenges, particularly in training methodologies and network optimization. Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[6]. Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9], have been developed to address these challenges, enabling the effective training of SNNs.
@@
  • Does bit plane representation enhance the performance of SNN?
  • What are impacts of different color models on SNN?
Especially, we observe that bit planes of an image contain pulse-like information, and can be applied to encode color information for SNNs to possibly improve accuracy.// After extensive experimentation, we concluded that bit planes contain valuable information that can be utilized in the rate coding process as a supplementary source to enhance the accuracy of direct surrogate backpropagation process [13] in SNN architectures. Furthermore, bit plane information proved to be reusable across multiple color models in most of the cases examined in our study.
A preliminary version of this paper is published in [14], where the benefit of including bit plane coding is proved by three grayscale datasets. In this extended version, the following new points have been added. First, the proposed method was updated to support color images. Second, eight more datasets are included. Third, impacts of color models are investigated.
@@ II Spiking Neural Network
In this section we present an overview of SNN knowledge related to our research.
II-A History of SNNs
Phenomenological neuron models can be progressively simplified until reaching a stage where even neural spikes are not modeled, leading to first and second generation neurons [15, 16]. This simplification significantly reduces computational demands but also decreases biological realism. The detail level in modeling often correlates with available computing power. For example, ANNs based on second generation neurons were historically limited by computing power [17], restricting the size of networks and the complexity of neuronal models. This limitation influenced the design of learning algorithms. Even with advancements in computing power, the complexity of neuronal models did not proportionately increase because existing learning algorithms were not compatible with more detailed models. Consequently, two distinct research fields emerged.
  • ANNs[18] focused on large networks of simplified neuron-like units, capable of learning through algorithms like backpropagation, and were successfully applied in pattern recognition, classification, and pattern completion tasks.
  • Computational Neuroscience[19] in contrast used detailed biophysical and phenomenological models in smaller networks to study electrophysiological processes, pattern generation, and dynamic behaviors of small neuron groups.
Although there have been attempts to create large networks of detailed biophysical neuron models [20], it remains challenging due to computational constraints.
Recent advances in computing power have blurred the lines between these fields. ANNs have become more biologically realistic, while networks of biophysical neurons have grown larger and more detailed. However, computational power still limits the application of detailed models in large neural networks for pattern recognition and classification. With improved computing power, new learning algorithms are being developed for these models. SNNs, which use spiking neurons modeled by phenomenological models like the Spike Response Model[21], represent the next step towards this goal. Due to the computational burden, biophysical models are less common in certain SNN applications. This paper focuses on research involving SNNs with phenomenological models, excluding those using biophysical models.
II-B The Integrate-and-Fire model
One of the earliest neuron models is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [22]. In this model, the neuron is characterized by its membrane potential $V$:
@@ II-B The Integrate-and-Fire model
While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons, challenges still remain in the conversion process. One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches. Despite the introduction of both data-driven and model-driven threshold determination methods[23], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts. Additionally, IF neurons lack time-dependent memory. Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires. This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [24]. Still, within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better LIF neurons[24], with the SEW-ResNet[25] architecture being one such example.
Figure 1: Illustrated bit planes of a sample image from CIFAR10[26] dataset.
II-C Deep Residual Learning in Spiking Neural Networks
@@
  • Inability to achieve identity mapping across neuron models: Residual learning is predicated on the concept of identity mapping, where layers can be trained to approximate an identity function. While this is straightforward in ResNets[27] with ReLU activation, it is not always feasible in Spiking ResNets[28]. Specifically, [25] claimed that for certain spiking neuron models, such as leaky integrate-and-fire (LIF) neurons[22] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging. This complicates the realization of identity mapping.
@@
SEW ResNet [25] was also the first work to propose and apply a surrogate gradient approach for optimizing SNNs. Using backpropagation methods to train SNNs [29, 13] has become a common practice in neuromorphic computing. However, applying backpropagation in SNNs presents certain challenges. In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
@@
As shown, directly differentiating $H(\theta,t)$ results in unstable gradients. To address this, [25] proposed substituting the derivative of the Heaviside function $\partial H$ with a numerically stable gradient from a different activation function during backpropagation, such as the derivative $\partial\sigma$ of the Sigmoid function $\sigma(\theta,t)$:
@@
This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively.
With these advantages, we decided to choose SEW Resnet[25] architecture for our study.
@@ III-B Color models in deep learning
Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks. For example, in [34], a research on traffic light recognition tested a range of color models across different deep learning architectures, with results indicating that the RGB model remains favored despite some of its limitations. This suggests that, while RGB is not perfect, it still outperforms other models in certain applications.
In another study on applying convolutional neural networks (CNNs) for skin detection against adversarial attacks of generative models[35], RGB and YCbCr color models demonstrated superior performance, particularly in handling varying lighting conditions. This highlights the adaptability of these color models in scenarios where lighting inconsistencies can degrade model accuracy.
Moreover, the development of ColorNet[11] illustrates the potential of leveraging multiple color models to capture a broader set of features. By combining information from these diverse color models, ColorNet outperforms traditional models that rely on a single color model. However, this approach introduces computational overhead due to the inclusion of smaller submodels, which can be a challenge despite the model’s increased efficiency in some cases.
Despite extensive researches on the use of color models in ANN, their application in SNN remains underexplored and requires further investigation.
@@ IV-A Hypothesis and Preliminary Study
To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[36], which is a simple dataset containing gray-scale images of hand-written numbers. For analysis in term of information contained in the coded spike signals, we perform the calculation of the mean base-2 Shannon entropy[37] $\overline{H}(X_{j})$ as:
$\overline{H}(X_{j})=-\frac{1}{N}\sum_{j=1}^{N}\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (5)
and the approximated average Kolmogorov complexity $\overline{K}$ across all dataset’s elements where the Kolmogorov complexity $K(X)$ of a coded image is estimated with DEFLATE compression algorithm[38] $C_{DEFLATE}(X)$ as:
$\overline{K}(X_{j})$ $=\frac{1}{N}\sum_{j=1}^{N}K(X_{j})$ (6)
$\approx\frac{1}{N}\sum_{j=1}^{N}C_{DEFLATE}(X_{j})$
$=\frac{1}{N}\sum_{j=1}^{N}Huffman(LZ77(X_{j}))$
where $Huffman(X)$ is the Huffman coding[39] and $LZ77(X)$ is the LZ77 compression[40] (or sometimes refered to as sliding window compression).
@@
We compared three approaches: traditional spike coding, bit plane coding only, and the combination of both spike coding and bit plane coding. The detailed settings of this experiment can be found in Section V-A. The results are provided in Table I.
Coding method Validation Accuracy (%) Averaged Kolmogorov complexity approximation (bits) Averaged Shannon entropy (bits)
Spike coding 98.62 721.5 0.53
Bit plane coding only 97.87 772.8 0.56
Combination of spike coding and bit plane coding 98.93 744.4 0.55
TABLE I: Average validation accuracy and analysis metrics comparison on MNIST[36]
@@
The results demonstrate that the approximated Kolmogorov complexity and Shannon entropy of bit-plane-coded data are higher, suggesting that the information content is more complex and less predictable. This may help explain why the accuracy of models using exclusively bit-plane-encoded inputs is lower compared to the baseline. On the other hand, a combination of rate coding and bit-plane coding resulted in an increase in performance. This outcome demonstrates that bit planes contain valuable information that SNN can learn from to improve their classification ability. This encourages us to carry out a further study and perform testing on other datasets.
IV-B Proposed Coding Method with Bit Planes
@@
As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding methodologies. Initially, the process commences with the conventional spike coding procedure applied to the image tensor. The bit plane coding algorithm (Algorithm 1) is employed to transform an input image into binary planes. The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from spike coding and bit plane coding. These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes.
Algorithm 1 is designed to encode an image tensor batch $X$ into $n_{bit}$ bit planes. To determine $n_{bit}$, it is essential to know the maximum possible value $X_{max}$ of the tensor batch. Since different color models have different $X_{max}$, this value must be predefined according to the selected color model. Once $X_{max}$ is defined, $n_{bit}$ can be calculate as $\left\lceil\log_{2}(X_{max})\right\rceil$.
To calculate the bit planes of an unnormalized integer input tensor batch $X$, we perform element-wise division of $X$ by 2. The remainder from this division constitutes a bit plane of $X$, while the quotient is saved for the extraction of subsequent bit planes. This process is repeated until we obtain $n_{bit}$ bit planes. The resulting bit planes are then concatenated to form $n_{bit}$ binary tensor batch of $X_{shape}$. This format closely resembles that of rate coding and ensures determinism in the coding results without having to sample more than $n_{bit}$ times.
Symbol Meaning
$X$ Input image tensors
$X_{max}$ Integer value indicate the maximum value of input image tensor
$log_{2}()$ Binary logarithm function
$X_{shape}$ Dimensions of X, usually include number of images being processed $B$, number of channels $C$ and resolution $H,W$
$\mathbb{Z}$ The set of all integers
TABLE II: Table of symbols used in Algorithm 1.
@@
function BitplaneEncode($X,X_{max}$)
@@
V Experiments
V-A Experimental settings

### v3

@@ Abstract
Spiking neural network (SNN) has emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint. However, their practical adoption is constrained by several challenges, prominently among them being performance optimization. In this study, we present a novel approach to enhance the SNN for image classification through a new coding method that exploits bit plane representation. Also, we investigate the impacts of color models of the proposed coding process. Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets of grayscale images and color images. To the best of our knowledge, this is the first research that considers bit planes and color models in the context of SNN. By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more effective SNNs models in future researches and applications. Source code for our experiments is available at https://github.com/luutn2002/bit-plane-snn
@@
Spiking neural network, spike coding, image classification, bit planes, color models
@@ I Introduction
SNNs utilize discrete spike events to transmit information between neurons, mirroring the asynchronous and event-driven communication observed in biological neurons[3]. This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[4]. The inherent sparsity and temporality in coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [5].
Despite their potential, SNNs pose significant challenges, particularly in training methodologies and network optimization. Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[6]. Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9, 10, 11], have been developed to address these challenges, enabling the effective training of SNNs.
@@
  • How can we utilize bit plane representation to enhance the performance of SNN?
  • How can we utilize different color models to enhance the performance of SNN?
In our research, we observe that bit planes of an image contain pulse-like information, and can be applied to encode pixel values for SNNs to improve accuracy during surrogate gradient training. Through extensive experiments, it is shown that bit planes contain valuable information that can be utilized in the spike coding process as a supplementary data source for SNN. Also, bit plane coding proves to be reutilized across multiple color models. Furthermore, RGB color model, which is the default color model of many datasets, is not always the best one among the evaluated color models in our study.
A preliminary version of this paper is published in [15], where the benefit of using bit plane coding is proved by three grayscale datasets. In this extended version, the following new points have been added.
  • First, the proposed method was updated to support color images.
  • Second, eight more datasets were included in the experiments.
  • Third, impacts of color models were investigated.
  • Fourth, discussion and analysis of the paper were significantly extended.
TABLE I: Table of abbreviation (in alphabetical order) used in this paper.
Abbreviation Full Name
ADD Addition
ANN Artificial neural network
CIE International Commission on Illumination
IF Integrate-and-fire
LIF Leaky integrate-and-fire
ReLU Rectified linear unit
SEW Spike-element-wise
SNN Spiking neural network
SNR Signal-to-noise ratio
STDP Spike-timing-dependent plasticity
TABLE II: Table of algorithmic symbols used in this paper.
Symbol Meaning
$X$ Input image tensors
$X_{max}$ Integer value indicate the maximum value of input image tensor
$log_{2}()$ Binary logarithm function
$X_{shape}$ Dimensions of X, include: $B$: number of images being processed. $C$: number of image channels. $H$: image height. $H$: image width.
$convertColor()$ Tensor color model converting function, convert $X$ to a specified color model decided by value of $colormodel$
$colormodel$ A string value specify the color model to convert the tensor to (i.e. ”RGB”, ”CMY”)
$mod$ Modulo division
$\mathbb{Z}$ The set of all integers
@@ II Spiking Neural Network
In this section we present an overview of SNN. The abbreviations and algorithmic symbols used throughout the paper are listed in Table I and Table II.
II-A Overview of SNNs
SNNs can be considered a biologically inspired variant of ANNs that offer a different computational paradigm. Unlike traditional ANNs, which process information in a continuous manner using floating-point activations and backpropagation-based learning [17], SNNs rely on discrete events called spikes to encode and transmit information [18]. This event-driven nature of SNNs allows for energy-efficient computations, making them suitable for neuromorphic hardware implementations [19]. While the majority of mainstream ANNs architecture operate on floating point data representations, SNNs mimic the sparse, time-dependent processing found in biological neurons [20]. Furthermore, SNNs employ temporal dynamics to learn and process temporal information naturally, leveraging mechanisms such as STDP for unsupervised learning [21]. Despite their advantages, the training of SNNs remains challenging due to the non-differentiable nature of spike events, often necessitating specialized training methods like surrogate gradients [9] or ANN-to-SNN conversion approaches [22]. Thus, SNNs present a compelling alternative to ANNs for energy-efficient and biologically plausible computations, while also introducing unique challenges in model design and training.
One of the earliest neuron models that is still widely applied in modern SNN architectures is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [23]. In this model, the neuron is characterized by its membrane potential $V$:
@@ II-A Overview of SNNs
While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons[24, 25, 26], challenges still remain in the conversion process. One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches. Despite the introduction of both data-driven and model-driven threshold determination methods[27], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts. Additionally, IF neurons lack time-dependent memory. Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires. This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [28].
Still, recent study have shown that SNN employed with IF neurons are more energy efficient than traditional ANN[29]. Within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better than leaky integrate-and-fire (LIF) neurons[28], with the SEW-ResNet[11] architecture being one such example.
Figure 1: Visualization of the extracted bit planes from a sample image in the CIFAR-10 dataset [30]. Higher-order bit planes (e.g., most significant bits) encode more spatial and structural information, capturing the prominent features of the image. In contrast, lower-order bit planes (e.g., least significant bits) primarily contain finer details and noise, contributing less to the overall image structure.
II-B Deep Residual Learning in SNNs
@@
  • Inability to achieve identity mapping across neuron models: Residual learning is predicated on the concept of identity mapping, where layers can be trained to approximate an identity function. While this is straightforward in ResNets[31] with ReLU activation, it is not always feasible in Spiking ResNets[26]. Specifically, [11] claimed that for certain spiking neuron models, such as LIF neurons[23] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging. This complicates the realization of identity mapping.
@@
In recent researches, using backpropagation methods to train SNNs [32, 10, 9, 11] has become a common practice in neuromorphic computing. In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
@@
As shown, directly differentiating $H(\theta,t)$ results in unstable gradients. To address this, we can substitute the derivative of the Heaviside function $\partial H$ with a numerically stable gradient from a different activation function during backpropagation, such as the derivative $\partial\sigma$ of the Sigmoid function $\sigma(\theta,t)$:
@@
This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively [11].
Similar to other new types of neural networks [33, 34, 35], the choice of input data representation may have strong impacts on output performance. Thus, in this paper we investigate a new way that employs bit plane coding for data representation in SNNs.
@@ III-A Bit planes in deep learning
In the context of digital discrete signals such as images, a bit plane represents the collection of bits that occupy the same position within the binary representation of the signal’s numerical values (as illustrated in Figure 1). For instance, in a 8-bit data representation, there are 8 distinct bit planes. The $1^{st}$ bit plane consists of the most significant bits, while the $8^{th}$ bit plane comprises the least significant bits. The first bit plane provides the coarsest yet most fundamental approximation of the signal’s values. As the bit plane index increases, its contribution to the overall representation diminishes. Consequently, incorporating additional bit planes progressively refines the signal’s approximation.
@@ III-B Color models in deep learning
Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks, such as image recognition[40, 13], image segmentation[14] and adversarial attacks prevention[41].
Despite extensive researches on the use of color models in ANN, their application in SNN remains unexplored and requires further investigation.
@@ IV Proposed Method
To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[42], which is a simple dataset containing grayscale images of hand-written numbers. We compared three approaches: traditional rate coding, bit plane coding only, and the combination of both rate coding and bit plane coding. The detailed settings of this experiment can be found in Section VI-A. The results are provided in Table III.
TABLE III: Average validation accuracy comparison on MNIST[42]
Coding method Validation Accuracy (%)
Rate coding 98.62
Bit plane coding only 97.87
Combination of rate coding and bit plane coding 98.93
@@
We can see that the accuracy when using exclusively bit-plane-encoded inputs is lower compared to the baseline. However, a combination of rate coding and bit-plane coding resulted in an increase in performance. This outcome demonstrates that bit planes contain valuable information that SNN can learn from to improve their classification ability. This encourages us to carry out a further study and perform testing on other datasets.
@@
As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding. Initially, the process commences with the conventional spike coding procedure applied to the image tensor. The bit plane coding algorithm is then employed to transform an input image into binary planes. The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from spike coding and bit plane coding. These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes.
The process of converting a multi-channel image into its bit planes involves isolating each bit of the binary representation of the pixel values for each color channel. For an $n_{bit}$ image, let $I_{C}(x,y)$ denote the intensity of a pixel at position $(x,y)$ for a given color channel $C$, pixel intensity can be expressed in terms of its binary representation as:
$I_{C}(x,y)=\sum_{k=0}^{n_{bit}}B_{k,C}(x,y)\cdot 2^{k}$ (5)
where $B_{k,C}(x,y)\in\{0,1\}$ is the $k$-th bit of the binary representation of $I(x,y)$, and $k$ represents the bit-plane index, with $k=0$ corresponding to the least significant bit and $k=n_{bit}$ corresponding to the most significant bit.
To isolate a specific bit-plane, the $k$-th bit of the pixel can be extracted using the following equation:
$B_{k,C}(x,y)=\left\lfloor\frac{I_{C}(x,y)}{2^{k}}\right\rfloor\mod 2$ (6)
@@
function BitplaneEncode($X,X_{max}$, $colormodel$)
     $X=convertColor(X,colormodel)$triangleright Convert $X$ to specified color model
@@
To encode an image tensor batch $X$ into $n_{bit}$ bit planes in our training process based on established knowledge, we implement Algorithm 1. It is essential to know the maximum possible value $X_{max}$ of the tensor batch in order to determine $n_{bit}$. Since different color models have different $X_{max}$, this value must be predefined according to the selected color model. Once $X_{max}$ is defined, $n_{bit}$ can be calculate as:
$n_{bit}=\lceil\log_{2}(X_{max})\rceil$ (7)
Bit planes calculation of an unnormalized integer input tensor batch $X$ can be performed by element-wise division of $X$ by 2. The remainder from this division constitutes a bit plane of $X$, while the quotient is saved for the extraction of subsequent bit planes. This process is repeated until we obtain $n_{bit}$ bit planes. The resulting bit planes are then concatenated to form $n_{bit}$ binary tensor batch of $X_{shape}$. This format closely resembles that of rate coding and ensures determinism in the coding results without having to sample more than $n_{bit}$ times.
V Theoretical basis
In this part, we aim to establish a theoretical foundation for our proposed method of spike coding within surrogate gradient approaches.
As discussed in Section II-B, surrogate gradient methods attempt to adapt deep learning’s gradient optimization techniques to SNNs, maintaining a close connection to traditional deep learning and convex optimization. A classical stochastic optimization problem associated with learning a target function $h$ can be formulated as[43]:
$\min_{\mathbf{w}}F_{h}(\mathbf{w})=\mathbb{E}_{\mathbf{x}}\left[\ell(p_{\mathbf{w}}(\mathbf{x}),h(\mathbf{x}))\right],$ (8)
where $\ell$ is a loss function, $\mathbf{x}$ represents stochastic inputs (assumed to be vectors in Euclidean space), and $p_{\mathbf{w}}$ is a differentiable predictor parameterized by $\mathbf{w}$. It is assumed that $E_{x}[\left\|\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}\right\|^{2}]\leq G(\mathbf{w})^{2}$ for some scalar $G(\mathbf{w})$, and that $F$ is differentiable[43, 44].
By reformulating our surrogate gradient optimization process as a traditional stochastic optimization problem, our objective can be defined as:
$\min_{\mathbf{w}}F_{h}(\mathbf{w})$ $=\mathbb{E}_{\mathbf{x}}\left[\ell(P_{\mathbf{w}}(\mathbf{x}),h(\mathbf{x}))\right],$ (9)
$P_{\mathbf{w}}(\mathbf{x})$ $=\mathbb{E}_{T}\left[p_{\mathbf{w}}(e(\mathbf{x}))\right],$
where the surrogate gradient model $P_{\mathbf{w}}(\mathbf{x})$ encodes the input $\mathbf{x}$ into a spike signal using an encoding function $e(\mathbf{x})$ and computes the expected output signal across the time dimension $T$ to generate the final stochastic prediction.
Previous studies in gradient-based optimization[43, 44] have demonstrated a positive correlation between a model’s performance and the signal-to-noise ratio (SNR) of the stochastic gradient, defined by:
$SNR(\delta(\mathbf{x}))=\frac{Sig(\delta(\mathbf{x}))}{Noi(\delta(\mathbf{x}))},$ (10)
$Sig(\delta(\mathbf{x}))=\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}$
$Noi(\delta(\mathbf{x}))=\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}$
$\delta(\mathbf{x})=h(\mathbf{x})g_{\mathbf{w}}(\mathbf{x})$
where $g_{{\mathbf{w}}}({\mathbf{x}})=\frac{\partial}{\partial{\mathbf{w}}}p_{{\mathbf{w}}}({\mathbf{x}})$ is the Jacobian of model $p_{{\mathbf{w}}}({\mathbf{x}})$ with respect to model weight $\mathbf{w}$. Since this correlation does not depend on input gradient calculation, we can directly apply this in our research.
To improve the SNR, we can either increase signal term $Sig(\delta(\mathbf{x}))$ or reduce noise term $Noi(\delta(\mathbf{x}))$. As shown in [11], the norm $\left\|\delta({\mathbf{x}})\right\|$ is positively correlated with timestep $T$:
$\left\|\delta({\mathbf{x}})\right\|=\left\|h^{\prime}({\mathbf{x}})\right\|\cdot\left\|\frac{\partial\ell}{\partial O^{l+k-1}}\right\|\cdot\left\|\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)\right\|$ (11)
$\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)=\sqrt{NT\alpha}$
$\alpha=\left[\Phi\,\Theta^{\prime}(1-V_{th})^{2k}+(1-\Phi)\,\Theta^{\prime}(0-V_{th})^{2k}\right]$
$\Phi=\frac{\sum_{j=0}^{N-1}\sum_{t=0}^{T-1}S^{l}_{j}[t]}{N\,T}$
where the equation denote the magnitude of $\left\|\delta({\mathbf{x}})\right\|$ with target function gradient $h^{\prime}({\mathbf{x}})$, previous blocks’ gradient $\frac{\partial L}{\partial O^{l+k-1}}$ of a $k$ sequential blocks $l$ layers Spiking-ResNet architecture[11] and Heaviside function’s gradient $\Theta^{\prime}$ while transmitting signal $S$ of previous layer output $O$ at firing rate $\Phi$ using $N$ neurons with threshold voltage $V_{th}$.
Upon performing signal concatenation using bit-plane-based signal $B$ of bit length $N_{bit}$ as in the proposed method, the magnitude of $\left\|\delta({\mathbf{x}})\right\|$ and firing rate $\Phi$ would increase as:
[…]
$\prod_{i=0}^{k-1}\Theta^{\prime}\left(S^{l+i}-V_{th}\right)\Theta^{\prime}\left(B^{l+i}-V_{th}\right)=\sqrt{N(T+N_{bit})\alpha}$
$\alpha=\left[\Phi\,\Theta^{\prime}(1-V_{th})^{2k}+(1-\Phi)\,\Theta^{\prime}(0-V_{th})^{2k}\right]$
$\Phi=\frac{\sum_{j=0}^{N-1}\sum_{t=0}^{T-1}S^{l}_{j}[t]+\sum_{j=0}^{N-1}\sum_{t=0}^{N_{bit}-1}B^{l}_{j}[t]}{N\,(T+N_{bit})}$
and improving the chance of obtaining higher $SNR(\delta(\mathbf{x}))$.
Figure 3: Comparing model gradient’s SNR of the proposed method (red), baseline (blue) and bit plane coding (green) of MNIST dataset across the epochs. Dashed line indicate the mean gradient SNR (higher is better based on our assumption).
To empirically validate our hypothesis, we measure the gradient SNR of the model across all training epochs on MNIST dataset (since to ensure that gradient does not saturate during training, $p_{\mathbf{w}}({\mathbf{x}})$ should serves as a reasonable approximation of the target function $h({\mathbf{x}})$). The resulted SNR measurements across all training epochs are presented in Figure 3. The findings demonstrate that our proposed method achieves a higher mean SNR during training process, which aligns with our initial assumption.
VI Experiments
VI-A Experimental settings